# Theoretical Detailed Analyses for DC readout and a Fabri-Pérot gravitational-wave detector

### v1

@@ V.3 Evaluation of Radiation pressure forces
Now, we evaluate the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega)D_{d}$. The details of the evaluation of these radiation pressure forces are given in Appendix B. In this section, we only explain the outline of Appendix B.
@@
To consider the situation where the state for the incident electric field from the light source in the coherent state with the complex amplitude $\alpha(\omega)$, we apply the displacement operator $D_{d}^{\dagger}$ from the left and $D_{d}$ from the right, which are defined by Eq. (101), to the obtained radiation pressure forces (326), (327), (329), and (331). Using (104), we separate the incident electric field quadrature $\hat{D}(\omega)$ from the light source into its classical part and the part of quantum fluctuations. Then, we evaluate the classical part, the linear-order contributions of the quantum fluctuations from $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$, and the linear-order contributions from the mirror displacements $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. After this evaluation, we consider the situation where the incident electric field from the light source is in the monochromatic coherent state with the central frequency $\omega_{0}$ through the choice of the complex amplitude $\alpha(\omega)$ for the coherent state by Eq. (108).
Thus, we reach the expressions of the Fourier transformation of the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ as Eqs. (338), (339), (340), and (341), respectively.
Here, we consider the tuning condition of the Fabri-Pérot cavity by
@@
Due to this tuning condition, the radiation pressure forces $D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$, $D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$, and $D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega)D_{d}$ given by Eqs. (338), (339), (340), and (341), respectively, are expressed as
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega)D_{d}$ (162)
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega)D_{d}$ (163)
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega)D_{d}$ (164)
@@
$D_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega)D_{d}$ (165)
@@ V.4 Solutions to the Heisenberg equations for mirrors’ motion
$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ (166)
@@
$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ (167)
@@
$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ (168)
@@
$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ (169)
@@
$D_{d}^{\dagger}\hat{Z}_{com}(\Omega)D_{d}$ $=$ $\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$ (172)
@@
$D_{d}^{\dagger}\hat{Z}_{dif}(\Omega)D_{d}$ $=$ $\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$ (173)
@@
$D_{d}^{\dagger}\hat{Z}_{comITM}(\Omega)D_{d}$ $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
@@
$D_{d}^{\dagger}\hat{Z}_{difITM}(\Omega)D_{d}$ $=$ $-\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right].$ (175)
We have to note that the classical constant terms in Eq. (172) and (174) diverge if we choose $\omega_{p}=0$, i.e., if we assume the mirrors are in completely free motion. This is the reason why we introduced the nonvanishing pendulum fundamental frequency $\omega_{p}$ in each equation of motion for mirrors. This introduction of the nonvanishing pendulum fundamental frequency $\omega_{p}$ is a natural consequence from the viewpoint of the actual ground-based gravitational-wave detectors. From a mathematical point of view, this introduction of $\omega_{p}$ is a kind of regularization of the divergence. However, we can regard the introduction of this regularization as reasonable for the actual ground-based gravitational-wave detectors. In this sense, we have to emphasize that this introduction $\omega_{p}$ has a physical meaning more than a mere mathematical regularization.
@@
VI Final input-output relation for a Fabri-Pérot Graviational-Wave Detector
@@
We also apply the approximation $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2\sim 0$ and $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2\sim 0$ as estimated in Eqs. (LABEL:eq:Omegatauy-taux-half) and (LABEL:eq:Omegatauy+taux-half). Furthermore, we consider the tuning condition of the Fabri-Pérot cavity by Eq. (161). In addition, we employ the tuning condition of the lengths between the BS and ITMs by
@@
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (178)
@@
$+i\frac{N\omega_{0}}{c}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right]$
$-i\frac{2NT\omega_{0}}{c}e^{\pm i\Omega\tau}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega)D_{d}+\cos\theta D_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega)D_{d}\right]$
$+\frac{iN\omega_{0}}{c}\left[i\sin\theta D_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega)D_{d}-\cos\theta D_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega)D_{d}\right].$
@@
$D_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega)D_{d}$ (179)
@@
Through the quadratures (180)–(184), variables $\mathfrak{R}$}}, $\beta$, $\kappa$, and $h_{SQL}$, the input-output relation (179) is given by
$D_{d}^{\dagger}\hat{b}_{\pm}(\Omega)D_{d}$ $=$ $+i\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\sin\theta\left[1+i{{\mathfrak{R}}}\right]2\pi\delta(\Omega)$ (189)
@@
$D_{d}^{\dagger}\hat{b}_{1}(\Omega)D_{d}$ $=$ $-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}{{\mathfrak{R}}}\sin\theta 2\pi\delta(\Omega)$ (193)
@@
$D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}$ $=$ $\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)$ (194)
@@
In Ref. [8], $D_{d}^{\dagger}\hat{b}_{2}D_{d}$ is regarded as the signal operator in gravitational-wave detectors. The first line of Eq. (194) is the contribution from the classical coherent light. The second line of Eq. (194) corresponds to the shot noise due to the errors of the photon number estimation in the coherent state of the laser. The third line of Eq. (194) is the contribution from the radiation pressure noise. Then, the final line of Eq. (194) corresponds to the gravitational-wave signal cooperating with the response function $e^{+i\beta}\sqrt{2\kappa}\cos\theta/h_{SQL}$.
@@
$\quad\times\left[\left(D_{d}^{\dagger}\hat{b}_{2}(\Omega)D_{d}-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)\right)\right$.
@@
without any reason except for the citation of Refs. [30, 31]. Incidentally, in Refs. [30, 31], there is no description that the noise spectral density $S_{(K)}(\Omega)$ defined by Eq. (196) is the actual quantum noise in measurements of gravitational waves. They only suggested that $S_{(K)}(\Omega)$ is an estimation measure of the amplitude of the quantum noises in the lasers of interferometers.
Besides the problem whether $S_{(K)}(\Omega)$ defined by Eq. (196) is an physically reasonable noise estimation or not, we can calculate $S_{(K)}(\Omega)$ through the definitions (196) of $S_{(K)}(\Omega)$ and the definition (195) of the noise operator $\hat{h}_{n}$. As a result of this calculation, we obtain the result
@@
where $\kappa^{(K)}$ and $h_{SQL}^{(K)}$ are defined Eqs. (206) and (209), respectively. This $S_{(K)}^{(K)}(\Omega)$ coincides with the noise spectral density for the Fabri-Pérot interferometer in Ref. [8].
@@
In the next section, we derive the quantum noise spectral density from the quantum fluctuations of photons at the photo-detector through a DC readout scheme, which is discussed in Sec. III.
VII DC readout scheme for a Fabri-Pérot Graviational-Wave Detector
Here, we apply the general arguments on the DC readout scheme for a Fabri-Pérot interferometer with the input-output relation (189) or equivalently Eqs. (193) and (194). Comparing Eqs. (19) and (20) and (189), we obtain

### v2

@@ V.3 Evaluation of Radiation pressure forces
Now, we evaluate the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\omega){{\mathfrak{D}}}_{d}$. The details of the evaluation of these radiation pressure forces are given in Appendix B. In this section, we only explain the outline of Appendix B.
@@
To consider the situation where the state for the incident electric field from the light source in the coherent state with the complex amplitude $\alpha(\omega)$, we apply the displacement operator ${{\mathfrak{D}}}_{d}^{\dagger}$ from the left and ${{\mathfrak{D}}}_{d}$ from the right, which are defined by Eq. (101), to the obtained radiation pressure forces (326), (327), (329), and (331). Using (104), we separate the incident electric field quadrature $\hat{D}(\omega)$ from the light source into its classical part and the part of quantum fluctuations. Then, we evaluate the classical part, the linear-order contributions of the quantum fluctuations from $\hat{D}_{v}(\omega)$ and $\hat{A}(\omega)$, and the linear-order contributions from the mirror displacements $\hat{X}_{x}$, $\hat{X}_{y}$, $\hat{X}_{XITM}$, and $\hat{X}_{YITM}$. After this evaluation, we consider the situation where the incident electric field from the light source is in the monochromatic coherent state with the central frequency $\omega_{0}$ through the choice of the complex amplitude $\alpha(\omega)$ for the coherent state by Eq. (108).
Thus, we reach the expressions of the Fourier transformation of the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ as Eqs. (338), (339), (340), and (341), respectively.
Here, we consider the tuning condition of the Fabry-Pérot cavity by
@@
Due to this tuning condition, the radiation pressure forces ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$, ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$, and ${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\Omega){{\mathfrak{D}}}_{d}$ given by Eqs. (338), (339), (340), and (341), respectively, are expressed as
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpXITM}(\Omega){{\mathfrak{D}}}_{d}$ (162)
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpYITM}(\Omega){{\mathfrak{D}}}_{d}$ (163)
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpx}(\Omega){{\mathfrak{D}}}_{d}$ (164)
@@
${{\mathfrak{D}}}_{d}^{\dagger}{{\mathscr{F}}}_{rpy}(\omega){{\mathfrak{D}}}_{d}$ (165)
@@ V.4 Solutions to the Heisenberg equations for mirrors’ motion
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ (166)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ (167)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ (168)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ (169)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\frac{N^{2}\hbar\omega_{0}}{c\omega_{p}^{2}}\left[1-\sqrt{1-T}\right]^{-2}\left[\frac{T}{m_{EM}}+\frac{2\left[\sqrt{1-T}-1+T\right]}{m_{ITM}}\right]2\pi\delta(\Omega)$ (172)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\frac{N\hbar\omega_{0}}{c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}\right]^{-1}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}$ (173)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\frac{2N^{2}\hbar\omega_{0}}{m_{ITM}c\omega_{p}^{2}}\sqrt{1-T}\left[1-\sqrt{1-T}\right]^{-1}2\pi\delta(\Omega)$
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\frac{2N\sqrt{1-T}\hbar\omega_{0}}{m_{ITM}c(\Omega^{2}-\omega_{p}^{2})}e^{+i\Omega\tau}\left[1-\sqrt{1-T}e^{+2i\Omega\tau}\right]^{-1}\cos(\Omega\tau)\left[\hat{A}(\Omega-\omega_{0})+\hat{A}(\Omega+\omega_{0})\right].$ (175)
We have to note that the classical constant terms in Eqs. (172) and (LABEL:eq:Eq-of-motion-XYITM-com-tuned-wobifreq-sol-sum-approx) diverge if we choose $\omega_{p}=0$, i.e., if we assume the mirrors are in completely free motion. This is the reason why we introduced the nonvanishing pendulum fundamental frequency, $\omega_{p}$, in each equation of motion for the mirrors. This introduction of the nonvanishing pendulum fundamental frequency $\omega_{p}$ is a natural consequence from the viewpoint of the actual ground-based gravitational-wave detectors. From a mathematical point of view, this introduction of $\omega_{p}$ is a kind of regularization of the divergence. However, we can regard the introduction of this regularization as reasonable for the actual ground-based gravitational-wave detectors. In this sense, we have to emphasize that this introduction of $\omega_{p}$ has a physical meaning more than a mere mathematical regularization.
@@
VI Final input-output relation for a Fabry-Pérot Graviational-Wave Detector
@@
In the DC readout scheme, we dare to set up the interferometer to leak the classical carrier field in the dark port to use this classical carrier field as a reference of the measurement of the small fluctuations in the dark port. Within this paper, we control this leakage of the classical carrier field through the control the parameter $\theta$ defined by Eq. (176), though there are many different methods to make a leakage of the classical carrier field at the dark port. We also apply the approximation $\Omega(\tau_{y}^{\prime}-\tau_{x}^{\prime})/2\sim 0$ and $\Omega(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2\sim 0$ as estimated in Eqs. (LABEL:eq:Omegatauy-taux-half) and (LABEL:eq:Omegatauy+taux-half). Furthermore, we consider the tuning condition of the Fabry-Pérot cavity by Eq. (161). In addition, we employ the tuning condition of the lengths between the BS and ITMs by
@@
Even if we specified the offset angle $\theta$ defined by Eq. (176), the average $(\tau_{x}^{\prime}+\tau_{y}^{\prime})/2$ is still free. Therefore, we have to tune this average distance to guarantee the output port is nearly dark.
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (178)
@@
$+i\frac{N\omega_{0}}{c}e^{\pm 2i\Omega\tau}\left[1-\sqrt{1-T}e^{\mp 2i\Omega\tau}\right]\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
$-i\frac{2NT\omega_{0}}{c}e^{\pm i\Omega\tau}\left[1-\sqrt{1-T}e^{\pm 2i\Omega\tau}\right]^{-1}\left[1-\sqrt{1-T}\right]^{-1}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{com}(\pm\Omega){{\mathfrak{D}}}_{d}+\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diff}(\pm\Omega){{\mathfrak{D}}}_{d}\right]$
$+\frac{iN\omega_{0}}{c}\left[i\sin\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{comITM}(\pm\Omega){{\mathfrak{D}}}_{d}-\cos\theta{{\mathfrak{D}}}_{d}^{\dagger}\hat{Z}_{diffITM}(\pm\Omega){{\mathfrak{D}}}_{d}\right].$
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{B}(\omega_{0}\pm\Omega){{\mathfrak{D}}}_{d}$ (179)
@@
Here, “SQL” in $h_{SQL}$ means “Standard Quantum Limit” which is given by Eq. (198) below. We may regard that the definition of $h_{SQL}$ comes from the inequality Eq. (198). Through the quadratures (180)–(184), variables $\mathfrak{R}$}}, $\beta$, $\kappa$, and $h_{SQL}$, the input-output relation (179) is given by
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{\pm}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $+i\sqrt{\frac{I_{0}}{\hbar\omega_{0}}}\sin\theta\left[1+i{{\mathfrak{R}}}\right]2\pi\delta(\Omega)$ (189)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{1}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}{{\mathfrak{R}}}\sin\theta 2\pi\delta(\Omega)$ (193)
@@
${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}(\Omega){{\mathfrak{D}}}_{d}$ $=$ $\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)$ (194)
@@
In Ref. [10], ${{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}{{\mathfrak{D}}}_{d}$ is regarded as the signal operator in gravitational-wave detectors. The first line of Eq. (194) is the contribution from the classical coherent light. The second line of Eq. (194) corresponds to the shot noise due to the uncertainties of the photon number estimation in the coherent state of the laser. The third line of Eq. (194) is the contribution from the radiation pressure noise. Then, the final line of Eq. (194) corresponds to the gravitational-wave signal cooperating with the response function $e^{+i\beta}\sqrt{2\kappa}\cos\theta/h_{SQL}$.
@@
$\quad\times\left[\left({{\mathfrak{D}}}_{d}^{\dagger}\hat{b}_{2}(\Omega){{\mathfrak{D}}}_{d}-\sqrt{\frac{2I_{0}}{\hbar\omega_{0}}}\sin\theta 2\pi\delta(\Omega)\right)\right$.
@@
without any reason except for the citation of Refs. [35, 36]. Incidentally, in Refs. [35, 36], there is no description that the noise spectral density $S_{(K)}(\Omega)$ defined by Eq. (196) is the actual quantum noise in measurements of gravitational waves. They only suggested that $S_{(K)}(\Omega)$ is an estimation measure of the amplitude of the quantum noises in the lasers of interferometers. However, the main purpose of Ref. [10] was to show some techniques to surpass the “standard quantum limit” using the squeezed states techniques and comparison with them, and it was not a serious problem of the definition of the noise spectral density at that time. After the publication of Ref. [10], general arguments of quantum measurement theory were proposed [14]. Then, the definitions and the meanings of “noises” began to be discussed seriously.
Besides the problem whether $S_{(K)}(\Omega)$ defined by Eq. (196) is a physically reasonable noise estimation or not, we can calculate $S_{(K)}(\Omega)$ through the definitions (196) of $S_{(K)}(\Omega)$ and the definition (195) of the noise operator $\hat{h}_{n}$. As a result of this calculation, we obtain the result
@@
where $\kappa^{(K)}$ and $h_{SQL}^{(K)}$ are defined Eqs. (206) and (209), respectively. This $S_{(K)}^{(K)}(\Omega)$ coincides with the noise spectral density for the Fabry-Pérot interferometer in Ref. [10].
@@
In the next section, we derive the quantum noise spectral density from the quantum fluctuations of photons at the photodetector through a DC readout scheme, which is discussed in Sec. III.
VII DC readout scheme for a Fabry-Pérot Graviational-Wave Detector
Here, we apply the general arguments on the DC readout scheme for a Fabry-Pérot interferometer with the input-output relation (189) or equivalently Eqs. (193) and (194). Comparing Eqs. (19) and (20) and (189), we obtain